CPU A runs at 1GHz and has average CPI of 3. CPU B runs at 2GHz and has average CPI of 4. When a benchmark program is compiled for CPU A it generates 500 instructions. When same program is compiled for CPU B, it generates 800 instructions. Which CPU is better. Why?
Answer #1

Let's calulate the execution time for both A and B.

Execution time = CPI * Clock Time * Instruction Count

_____________________________________

For A :

CPI = 3, Clock time = 1 / Clock rate = 1 / 109 = 1 ns

Instruction Count = 500

Execution time = 3 * 1 * 500 = 1500 ns

______________________________________

For B :

CPI = 4, Clock time = 1 / Clock rate = 1 / 2 * 109 = 0.5 ns

Instruction Count = 800

Execution time = 4 * 0.5 * 800 = 1600 ns

_______________________________________

So, CPU A has less execution time so it means CPU A is faster.

Thus, CPU A is better.


answered by: ANURANJAN SARSAM
Answer #2

Let's calulate the execution time for both A and B.

Execution time = CPI * Clock Time * Instruction Count

_____________________________________

For A :

CPI = 3, Clock time = 1 / Clock rate = 1 / 109 = 1 ns

Instruction Count = 500

Execution time = 3 * 1 * 500 = 1500 ns

______________________________________

For B :

CPI = 4, Clock time = 1 / Clock rate = 1 / 2 * 109 = 0.5 ns

Instruction Count = 800

Execution time = 4 * 0.5 * 800 = 1600 ns

_______________________________________

So, CPU A has less execution time so it means CPU A is faster.

Thus, CPU A is better.

_________________________________________

Let me know if you have any doubts in the comments. Please upvote if the answer helped you.